Garrett Corp. has been going through a difficult financial period. Over the past three years, it's stock price has dropped from $50 to $18 per share. Throughout this downturn, Garrett has managed to pay a $1 dividend each year. Management feels the worst is over but intends to maintain the $1 dividend for three more years, after which they plan to increase it by 6% per year indefinitely. Comparable stocks are returning at 11%.

a. If these projections are accurate, is Garrett stock a good buy at $18?

b. How do you think the market feels about Garrett's management?
Answer #1

Horizon value = P3 = D4/(k-g)

= 1*106%/(11%-6%)

= 1.06/5%

= 21.2

Current price = D1/(1+k)+ D2/(1+k)^2 + (D3+P3) /(1+k)^3

= 1/1.11^1+1/1.11^2+(1+21.2) /1.11^3

= 17.94

a: No, since the intrinsic value is lesser than the market price.

b: The management is lieberal with its dividend policy. The management is not retaining profits even in low times.